Rita has a collection of 105 coins consisting of nickels, dimes, and quarters. The number of dimes is 5 more than one-third the number of nickels, and the number of quarters is twice the number of dimes. How many coins of each kind does she have?
Answer by stanbon(75887)

Rita has a collection of 105 coins consisting of nickels, dimes, and quarters. The number of dimes is 5 more than one-third the number of nickels, and the number of quarters is twice the number of dimes. How many coins of each kind does she have?
------
Quantity Equation: n + d + q = 105 coins
Quantity Equation: d = (1/3)n + 5
Quantity Equation: q = 2d
--------
Substitute for d and q and solve for "n":
n + (1/3)n + 5 + 2d = 105
------
(4/3)n + 5 + (2/3)n + 10 = 105
2n + 15 = 105
------
2n = 90
n = 45 (number of nickels)
d = (1/3)n+5 = 15+5 = 20 (number of dimes)
q = 2d = 40 (number of quarters)
==================================
Cheers,
Stan H.